the perimeter of a rectangle whose length is 3 less than five times its width

Answer by checkley79(3341)

P=2L+2W
L=5W-3
P=2(5W-3)+2W
P=10W-6+2W
P=12W-6
12W=P+6
W=(P+6)/12 ANS. FOR THE WIDTH.